# Hot Rocks Survey III: A deep eclipse for LHS 1140c and a new Gaussian process method to account for correlated noise in individual pixels

### v1

@@ 5 Modelling
where is the stellar effective temperature, is the stellar radius, is the semi-major axis, is the surface albedo (i.e. the Bond albedo of the surface), is the surface emissivity and is the heat redistribution factor, where represents full heat redistribution while represents zero heat redistribution (Hansen 2008). This equation is adapted from Malik et al. (2019a) and assumes there is no extra heat emitted due to a high interior planetary temperature. For the bare rock case we set as we expect insignificant heat redistribution without an atmosphere (Joshi et al. 1997; Selsis et al. 2011; Koll 2022). For all simulations, we assumed that the surface albedo was constant in wavelength which implies (as emission and absorption can occur at different wavelengths this is not necessarily true otherwise). Emission was then calculated using Planck’s law with the dayside temperature and scaling by the emissivity .
We also added the reflected light component:
@@
For the atmospheric cases, we used the 1D atmospheric model HELIOS (Malik et al. 2017, 2019b) with a solid surface bottom layer (Malik et al. 2019a; Whittaker et al. 2022) to calculate the planet’s temperature structure and its synthetic emission and reflection spectra. The temperature profile of a planet’s atmosphere is primarily determined by the balance between radiative, convective processes and heat redistribution. The model considers well-mixed atmospheres with various gas compositions consisting of CO2, H2O, SO2, N2 and O2. It employs k-distribution tables for opacities, integrating radiative fluxes over 386 spectral bands and 20 Gaussian points. Opacities are calculated using the HELIOS-K (Grimm & Heng 2015) code and using cross-sections for CO2 from Rothman et al. (2010), H2O from Barber et al. (2006), SO2 from Underwood et al. (2016) and O2 from Gordon et al. (2022). Additionally, we incorporate Rayleigh scattering cross-sections for H2O, CO2, N2 and O2 (Cox 2000; Sneep & Ubachs 2005; Thalman et al. 2014), but did not include scattering for SO2. For all atmospheric models, the surface albedo was fixed to .
@@
Figure 15 shows a range of models including pure CO2, pure H2O, pure SO2 and a CO2/H2O mixture. Each model was run 100 times with the input planetary parameters , , (planetary surface gravity) and stellar parameters , , randomly drawn from their constraints listed in Cadieux et al. (2024a) to propagate the uncertainty in these parameters. The figure displays the emission spectra resulting from the modelling, with the shaded regions representing the 1 range of the computed values.
We find that the eclipse depths from aperture photometry, pixel-fitting and the independent analysis using optimal extraction are all in excellent agreement with a low albedo bare rock (). Higher albedo surfaces of diverge at (uncertainties given relative to the joint-fit using aperture extraction with GPs; ppm). Extremely tenuous (¡1 mbar) CO2 or H2O atmospheres naturally behave similarly to bare rocks and cannot be distinguished using these observations.
@@
As we may expect volcanically outgassed secondary atmospheres to include sulphur-based molecules such as SO2, we also tested a range of pure SO2 atmospheres. In reality, sulphur chemistry is complex and SO2 is very photochemically sensitive, so more sophisticated photochemical modelling would need to be performed to understand what the dominant form of sulphur would be for this planet (see e.g. Bello-Arufe et al. 2025). It is likely that a dense SO2 atmosphere would require constant replenishment from volcanic outgassing, which given the low eccentricity of the planet ( to 95% confidence, see Appendix C) would not be expected from tidal forces, unlike suggestions for the L 98-59 system (Seligman et al. 2024). Noting these caveats, we recover that pure 1 bar SO2 atmospheres are ruled out at when modelling heat redistribution (plotted in Figure 15), while without heat redistribution this result decreases to . Atmospheres of 100 mbar or weaker are all within regardless of the effect of heat redistribution. We note that longer wavelengths such as those covered by the F1800W filter might provide significantly tighter constraints for thinner SO2 atmospheres. We have included models of SO2 in a CO2 background in Appendix E which suggest that combinations of these molecules may produce particularly strong signatures which we can rule out down to 1 mbar. However, we again note that these models exclude any photochemical or geochemical modelling and we hope future work can examine to what significance physically plausible volcanically outgassed atmospheres are ruled out to.
@@
0.1 mbar | 0.6 | 0.6 | 0.6 |
1 mbar | 0.6 | 0.6 | 0.7 |
5 mbar | 0.6 | 0.7 | 0.8 |
10 mbar | 0.6 | 0.7 | 0.9 |
50 mbar | 0.6 | 1.0 | 1.9 |
100 mbar | 0.7 | 1.5 | 2.7 |
500 mbar | 0.9 | 3.1 | 4.0 |
1 bar | 1.2 | 3.7 | 4.3 |
5 bar | 2.7 | 4.5 | 4.8 |
10 bar | 3.4 | 4.8 | 5.0 |
@@ 6 Discussion
These constraints are very informative because the planet’s low equilibrium temperature of K (Cadieux et al. 2024a), its relatively mild present-day radiation environment (Spinelli et al. 2019) and the reported age of the system (¿5 Gyr; Dittmann et al. 2017), make it one of the better known candidates to have an outgassed atmosphere or to have retained some of its primary atmosphere (Kite & Barnett 2020; Chatterjee & Pierrehumbert 2024). In contrast, our result is in line with the suggestion that historic levels of XUV fluence may be underestimated for mid-type M dwarfs like LHS 1140, making atmospheric retention more difficult (Pass et al. 2025). As this star also hosts LHS 1140b - a larger Super-Earth which likely still retains an atmosphere555There is still some discussion on the composition of LHS 1140b, see Cadieux et al. (2024b) and Damiano et al. (2024) for the evidence distinguishing rocky, water world or cloudy mini-Neptune scenarios. - our new atmospheric constraints should help make the LHS 1140 system an important benchmark for understanding atmospheric escape and the cosmic shoreline (Zahnle & Catling 2017).
We have introduced a new technique which can act as an alternative to aperture photometry or optimal extraction. Section 3 demonstrated that our method presents a clear advantage when we have individual pixels contaminated by systematics, as it can weight away from them towards cleaner pixels. For the JWST observations, we saw in our second eclipse that the pixel-fitting method could weight away from two discrepant pixels apparently affected by a persistence effect from a cosmic ray. While the method has disadvantages, such as its increased computational cost and complexity to implement, it provides a new way to analyse and explore data at the pixel-level which can identify problematic pixels in an empirical way. It’s also important to note that while our observations do not appear heavily impacted by systematics, an observation of another target in our survey - LHS 1478b - appears to be significantly impacted by the presence of systematics (August et al. 2025) which our method may be able to alleviate. The fact that our method is capable of distinguishing between systematics in individual pixels or across all pixels could also help to diagnose the cause of these systematics.
@@
The pixel-fitting method has huge potential as it could be applied to a wide range of existing time-series observations including spectroscopic time-series. There is currently a move within exoplanet atmospheres research to analyse JWST spectroscopic time-series at the native/pixel resolution, yet there is concern over how much correlation this introduces in the noise between neighbouring light curves, with standard analysis approaches typically ignoring correlations between light curves (see Ih & Kempton 2021 and Fortune et al. 2024 for how this can affect atmospheric retrievals). Note that MIRI/LRS uses the same detector as MIRI/Imaging (but on a subarray different to our observations), so many of the detector systematics seen in our observations - such as the flux-conserving systematics - might affect MIRI/LRS observations too. Future work extending pixel-fitting to spectroscopy could examine whether this anti-correlation between neighbouring pixel light curves could explain the large increase in scatter in transmission spectra noted when extracting spectroscopic light curves with one pixel wide bins (Bell et al. 2024). Implementing pixel-fitting for spectroscopy would differ as the eclipse/transit depth would not be shared across all pixels but only between pixels at the same wavelength. Computational cost could be a barrier to implementation as the spectral trace is spread across more pixels. Strategies such as binning in time and parallelising over multiple GPUs could help to mitigate this however.
Finally, we believe the relationship between the initial detector settling slope and the previous filter used by MIRI warrants further study as a significant fraction of MIRI TSOs are affected by this and mitigating it could provide a serious improvement to the efficiency of the instrument. A strategy such as changing when the filter is moved into place and/or changing how MIRI is clocked in idle mode could have a large effect on the initial settling slope.
@@ 7 Conclusions
Our three eclipses of LHS 1140c using MIRI/Imaging at 15m have allowed us to rule out a wide range of possible atmospheres. We have found that LHS 1140c is consistent with a low albedo bare rock and is unlikely to host an atmosphere capable of substantial heat redistribution, or an atmosphere containing significant levels (100 ppm) of CO2. Using an absolute flux calibration, we have measured its dayside brightness temperature to be K, close to the maximum expected dayside temperature of K. These results are fully consistent with LHS 1140c lacking any substantial atmosphere. There are alternative possibilities, including dense atmospheres which are very optically thin, but these models may require some fine tuning to explain the lack of strong absorbers like CO2. If LHS 1140c does lack an atmosphere then it could have strong implications for the cosmic shoreline model of atmospheric escape, as LHS 1140c is a super-Earth with one of the lowest instellations of any target this technique has been applied to so far (after TRAPPIST-1b and c; Greene et al. 2023; Zieba et al. 2023b). Given that the outer planet LHS 1140b is also a super-Earth but is likely to host an atmosphere (Damiano et al. 2024; Cadieux et al. 2024b), this could place each planet on either side of the cosmic shoreline.
We have identified a possible relationship in the strength and sign of the initial settling slope with the previous filter used by MIRI. This may help to inform ways to mitigate this effect in future.
We have also demonstrated a new method of analysing time-series photometry which produces similar results to aperture photometry on these datasets. The benefits of this method are:
It can identify and weight away from systematics arising in individual pixels, which can lead to more precise and robust measurements.
It can identify if systematics arise from individual pixels or are common to all pixels, helping to determine the source of the systematics.
It can weight towards higher signal-to-noise pixels similar to optimal extraction.
It retrieves additional information about the detector such as how the PSF behaves in time and the strength of interpixel capacitance.
It can constrain information about the amplitude of background scatter remaining after the background subtraction.
@@ Appendix A Cosmic ray persistence effect
In order to verify that the persistence effect seen in Figure 1 was caused by a cosmic ray, we took the uncalibrated ‘uncal’ files for the second eclipse observation and calculated the change in data number between the groups before and after the cosmic ray hit. This occurred between groups 7 and 8 of the 524th integration. To correct for the expected increase in data number due to flux from the target star, thermal background, etc. we subtracted the median frame of all integrations divided by the number of groups. This procedure had minimal effect as this expected increase in data number was far smaller (of the order of hundreds of data numbers) than the size of the jump from the cosmic ray (16,000 data numbers). In Figure 17 we overlay this increase in data number between these two groups with the pixel light curves centred around this cosmic ray event. These pixel light curves are from the reduction described in Section 4.1 which used the jump step correction and also performed an additional custom median outlier clipping step on the stage 1 and 2 processed data. These steps remove the significant jump in flux observed in many of these light curves in the integration containing the cosmic ray hit, but they do not perform any correction for this longer duration persistence artifact.
@@
In summary, persistence effects due to cosmic rays hitting individual pixels may be a widespread phenomenon for which the pixel-fitting method could be beneficial in order to weight away from. If these persistence effects could be modelled then it may be possible to directly fit them in the mean function of the pixel-fitting method (e.g. fitting an exponential ramp to the affected pixels) rather than weight away from them, potentially improving signal-to-noise. This could be more important if a strong cosmic ray were to affect one of the central pixels of the PSF or in the case of a high rate of cosmic ray hits as in the eclipse of GJ-357b.
@@
Appendix B Effect of intra-pixel sensitivity variations and flat fielding errors
@@
As described in Section 2.1, there are multiple reasons why intrapixel sensitivity should be significantly reduced in MIRI time-series compared to Spitzer. To confirm that it can be assumed to be negligble for MIRI, we performed tests to examine to what level intra-pixel sensitivity could affect our observations. A related issue is flat fielding, while we did run the flatfield step in the JWST pipeline for the pixel-fitting analyses, any inaccuracies in the flat fielding would result in uncorrected for variations in sensitivity between different pixels (in contrast to intra-pixel sensitivity which describes sensitivity variations within a pixel). We can examine both of these effects using simulated observations, similar to what has been performed for Spitzer (Morello 2015).
@@
All of our tests follow the same procedure. First, we generate a possible pixel sensitivity map. As we were unable to find any measurements of intrapixel sensitivity for MIRI in the literature, we tested various possibilities, shown in Figure 18. We considered a ”gradient” map in which the pixel surface smoothly varies in sensitivity with a constant gradient from a relative efficiency of 95% in the top left corner to 105% in the bottom right corner. We also considered a ”quadratic” map where the efficiency varied as a function of the square of the distance from the centre of the pixel. The normalised map has an efficiency ranging from 103% at the pixel centre to 93% at the edges. Our ”random” sensitivity map was generated using a Gaussian process with a 2D squared-exponential kernel function with correlation a function of the and position within a pixel. The height scale for this GP produced sensitivity variations between 95% to 105% relative efficiency. Each map was generated as a grid of points within each pixel. We then copied this map for a grid of pixels, creating an overall sub-pixel resolution sensitivity map simulating 225 total pixels.
@@
In addition to testing the scenario where all pixels share the same sensitivity map, we also generated maps which used randomised gradient maps, quadratic maps and random 2D Gaussian process maps which were independent for each pixel. These maps also exhibit 10% variations in efficiency across each pixel. To test flat fielding errors, we generated sub-pixel maps which had constant efficiency across each surface, but we varied this constant for each pixel with a standard deviation of 1% and centred around 100%. The MIRI reference file for flat fielding exhibits 1% level variations between pixels, so if we run the flatfield step then we are modelling these maps being incorrect with a standard deviation of 1%.
@@
Our results were that all of the intra-pixel sensitivity maps and flat fielding errors were completely negligible, never producing greater than a couple ppm of additional noise in our aperture extracted light curves (see Table 9). The pointing variation shows minimal correlation in time so most of this variation behaves like white noise and is tiny compared to the level of photometric noise in our observations (800 ppm). Note that the noise would be added in quadrature to photon noise, so even 40 ppm of noise from intra-pixel sensitivity would only increase the total aperture extracted noise by ppm in combination with 800 ppm of photon noise. These effects are therefore more than an order of magnitude below the level that would impact our observations, meaning that significantly more extreme sub-pixel sensitivity variations would be needed to have any noticeable effect. While the maximum effect in a particular pixel could be a couple hundred ppm (see the ”Max pixel (ppm)” column in Table 9), this was largely only in the outer regions of the PSF with the lowest signal-to-noise, and the photon noise in any individual pixel for the real observations is never better than 3000 ppm even in the central pixels, so this effect should also be negligible for the pixel-fitting method. The exact values in Table 9 would vary if different random draws of sensitivity maps were generated, but as the effect is so far below the level of photon noise it appears unlikely this would ever have a significant impact.
It is worth repeating that the Si:As detector technology used for MIRI is expected to show very minor intra-pixel sensitivity, so extreme variations across each pixel surface is expected to be highly unlikely. This is similar to what was found with the Si:As detectors on Spitzer, which did not display the same level of intrapixel sensitivity as the shorter wavelength m and m channels which used InSb detectors (Knutson et al. 2008). The narrower PSF widths from the shorter wavelength filters (e.g. F560W) may show slightly more variation, although rerunning the same tests with a reduced FWHM = 1.882px matching that of the F560W filter (and setting the aperture radius to match this FWHM), we still find that all maps produce ppm of noise. This demonstrates that the exquisite pointing stability of JWST is enough to result in minimal intrapixel sensitivity, even if the pixels do demonstrate 10% level variations in sensitivity across their surfaces. Therefore, it appears unlikely that intra-pixel sensitivity or flat fielding errors have any significant effect on time-series observations with MIRI/Imaging.

### v2

@@ 5 Modelling
where is the stellar effective temperature, is the stellar radius, is the semi-major axis, is the surface albedo (i.e. the Bond albedo of the surface), is the surface emissivity, and is the heat redistribution factor, where represents full heat redistribution, while represents zero heat redistribution (Hansen 2008). This equation is adapted from Malik et al. (2019a) and assumes there is no extra heat emitted due to a high interior planetary temperature. For the bare rock case we set as we expect insignificant heat redistribution without an atmosphere (Joshi et al. 1997; Selsis et al. 2011; Koll 2022). For all simulations, we assumed that the surface albedo was constant in wavelength, which implies (as emission and absorption can occur at different wavelengths this is not necessarily true otherwise). Emission was then calculated using Planck’s law with the dayside temperature and scaling by the emissivity .
We also added the reflected light component,
@@
For the atmospheric cases, we used the 1D atmospheric model HELIOS (Malik et al. 2017, 2019b) with a solid surface bottom layer (Malik et al. 2019a; Whittaker et al. 2022) to calculate the planet’s temperature structure and its synthetic emission and reflection spectra. The temperature profile of a planet’s atmosphere is primarily determined by the balance between radiative, convective processes and heat redistribution. The model considers well-mixed atmospheres with various gas compositions consisting of CO2, H2O, SO2, N2 and O2. It employs k-distribution tables for opacities, integrating radiative fluxes over 386 spectral bands and 20 Gaussian points. Opacities are calculated using the HELIOS-K (Grimm & Heng 2015) code and using cross-sections for CO2 from Rothman et al. (2010), H2O from Barber et al. (2006), SO2 from Underwood et al. (2016) and O2 from Gordon et al. (2022). Additionally, we incorporate Rayleigh scattering cross-sections for H2O, CO2, N2, and O2 (Cox 2000; Sneep & Ubachs 2005; Thalman et al. 2014), but did not include scattering for SO2. For all atmospheric models, the surface albedo was fixed to .
@@
Figure 15 shows a range of models including pure CO2, pure H2O, pure SO2, and a CO2/H2O mixture. Each model was run 100 times with the input planetary parameters , , (planetary surface gravity) and stellar parameters , , randomly drawn from their constraints listed in Cadieux et al. (2024a) to propagate the uncertainty in these parameters. The figure displays the emission spectra resulting from the modelling, with the shaded regions representing the 1 range of the computed values.
We find that the eclipse depths from aperture photometry, pixel-fitting and the independent analysis using optimal extraction are all in excellent agreement with a low-albedo bare rock (). Higher albedo surfaces of diverge at (uncertainties given relative to the joint-fit using aperture extraction with GPs; ppm). Extremely tenuous (¡1 mbar) CO2 or H2O atmospheres naturally behave similarly to bare rocks and cannot be distinguished using these observations.
@@
As we may expect volcanically outgassed secondary atmospheres to include sulphur-based molecules such as SO2, we also tested a range of pure SO2 atmospheres. In reality, sulphur chemistry is complex and SO2 is very photochemically sensitive; thus, more sophisticated photochemical modelling would need to be performed to understand what the dominant form of sulphur would be for this planet (see e.g. Bello-Arufe et al. 2025). It is likely that a dense SO2 atmosphere would require constant replenishment from volcanic outgassing, which given the low eccentricity of the planet ( to 95% confidence, see Appendix C) would not be expected from tidal forces, unlike suggestions for the L 98-59 system (Seligman et al. 2024). Noting these caveats, we recover that pure 1 bar SO2 atmospheres are ruled out at when modelling heat redistribution (plotted in Figure 15); however, without any heat redistribution, this result would decrease to . Atmospheres of 100 mbar or weaker are all within regardless of the effect of heat redistribution. We note that longer wavelengths such as those covered by the F1800W filter might provide significantly tighter constraints for thinner SO2 atmospheres. We have included models of SO2 in a CO2 background in Appendix E which suggest that combinations of these molecules may produce particularly strong signatures which we can rule out down to 1 mbar. However, we again note that these models exclude any photochemical or geochemical modelling and we hope future work can examine to what significance physically plausible volcanically outgassed atmospheres are ruled out to.
@@
| 0.1 mbar | 0.6 | 0.6 | 0.6 |
| 1 mbar | 0.6 | 0.6 | 0.7 |
| 5 mbar | 0.6 | 0.7 | 0.8 |
| 10 mbar | 0.6 | 0.7 | 0.9 |
| 50 mbar | 0.6 | 1.0 | 1.9 |
| 100 mbar | 0.7 | 1.5 | 2.7 |
| 500 mbar | 0.9 | 3.1 | 4.0 |
| 1 bar | 1.2 | 3.7 | 4.3 |
| 5 bar | 2.7 | 4.5 | 4.8 |
| 10 bar | 3.4 | 4.8 | 5.0 |
@@ 6 Discussion
These constraints are very informative because the planet’s low equilibrium temperature of K (Cadieux et al. 2024a), its relatively mild present-day radiation environment (Spinelli et al. 2019), and the reported age of the system (¿5 Gyr; Dittmann et al. 2017), make it one of the better known candidates to have an outgassed atmosphere or to have retained some of its primary atmosphere (Kite & Barnett 2020; Chatterjee & Pierrehumbert 2024). In contrast, our result is in line with the suggestion that historic levels of XUV fluence may be underestimated for mid-type M dwarfs such as LHS 1140, making atmospheric retention more difficult (Pass et al. 2025). As this star also hosts LHS 1140b – a larger Super-Earth which likely still retains an atmosphere888There is still some discussion on the composition of LHS 1140b, see Cadieux et al. (2024b) and Damiano et al. (2024) for evidence distinguishing rocky, water world, or cloudy mini-Neptune scenarios. – our new atmospheric constraints are expected to help make the LHS 1140 system an important benchmark for understanding atmospheric escape and the cosmic shoreline (Zahnle & Catling 2017).
We have introduced a new technique, which can act as an alternative to aperture photometry or optimal extraction. Section 3 demonstrates that our method presents a clear advantage when we have individual pixels contaminated by systematics, as it can weight away from them towards cleaner pixels. For the JWST observations, we see in the case of our second eclipse that the pixel-fitting method could weight away from two discrepant pixels apparently affected by a persistence effect from a cosmic ray. While the method has disadvantages, such as its increased computational cost and complexity to implement, it provides a new way to analyse and explore data at the pixel-level which can identify problematic pixels in an empirical way. It is also important to note that while our observations do not appear heavily impacted by systematics, an observation of another target in our survey (LHS 1478b) appears to be significantly impacted by the presence of systematics (August et al. 2025), which our method may be able to alleviate. The fact that our method is capable of distinguishing between systematics in individual pixels or across all pixels could also help to diagnose the cause of these systematics.
@@
The pixel-fitting method has huge potential as it could be applied to a wide range of existing time-series observations including spectroscopic time series. There is currently a move within exoplanet atmospheres research to analyse JWST spectroscopic time series at the native or pixel resolution, yet there is concern over how much correlation this introduces in the noise between neighbouring light curves, with standard analysis approaches typically ignoring correlations between light curves (see Ih & Kempton 2021 and Fortune et al. 2024 for how this can affect atmospheric retrievals). We note that MIRI/LRS uses the same detector as MIRI/Imaging (but on a subarray that is different from that of our observations); thus, many of the detector systematics seen in our observations (e.g. the flux-conserving systematics) might affect MIRI/LRS observations too. Future work extending pixel-fitting to spectroscopy could examine whether this anti-correlation between neighbouring pixel light curves could explain the large increase in scatter in transmission spectra noted when extracting spectroscopic light curves with one pixel-wide bins (Bell et al. 2024). Implementing pixel-fitting for spectroscopy would differ as the eclipse depth would not be shared across all pixels – but only between pixels at the same wavelength. The computational cost could be a barrier to implementation as the spectral trace is spread across more pixels. Strategies such as binning in time and parallelising over multiple GPUs could help to mitigate this.
Finally, we believe the relationship between the initial detector settling slope and the previous filter used by MIRI warrants further study as a significant fraction of MIRI TSOs are affected by this. Therefore, mitigating these effects could provide a serious improvement to the efficiency of the instrument. A strategy such as changing when the filter is moved into place and/or changing how MIRI is clocked in idle mode could have a large effect on the initial settling slope.
@@ 7 Conclusions
Our three eclipses of LHS 1140c using MIRI/Imaging at 15m have allowed us to rule out a wide range of possible atmospheres. We have found that LHS 1140c is consistent with a low-albedo bare rock and is unlikely to host an atmosphere capable of substantial heat redistribution or an atmosphere containing significant levels (100 ppm) of CO2. Using an absolute flux calibration, we measured its dayside brightness temperature to be K, close to the maximum expected dayside temperature of K. These results are fully consistent with the scenario that LHS 1140c lacks a substantial atmosphere. There are alternative possibilities, including dense atmospheres that are very optically thin, but these models may require some fine-tuning to explain the lack of strong absorbers such as CO2. If LHS 1140c does indeed lack an atmosphere, then it could have strong implications for the cosmic shoreline model of atmospheric escape, as LHS 1140c is a super-Earth with one of the lowest instellations of any target this technique has been applied to so far (after TRAPPIST-1b and c; Greene et al. 2023; Zieba et al. 2023b). Given that the outer planet LHS 1140b is also a super-Earth, but one that is likely to host an atmosphere (Damiano et al. 2024; Cadieux et al. 2024b), this could place each planet on either side of the cosmic shoreline.
We have identified a possible relationship in the strength and sign of the initial settling slope with the previous filter used by MIRI. This may help to inform ways to mitigate this effect in future. We have also demonstrated a new method of analysing time series photometry, which produces similar results to aperture photometry on these datasets. The benefits of this method include:
identifying and weighting away from systematics arising in individual pixels, potentially leading to more precise and robust measurements.
identifying whether systematics arise from individual pixels or are common to all pixels, helping to determine the source of the systematics.
weighting towards higher S/N pixels, similar to optimal extraction.
retrieving additional information about the detector, such as how the PSF behaves over time and the strength of interpixel capacitance.
constraining information about the amplitude of background scatter remaining after the background subtraction.
Overall, we hope this paper provides greater insight into the systematics present in MIRI data and provides useful approaches for dealing with them. This may help to robustly characterise more exoplanets in the future.
@@ Appendix A Cosmic ray persistence effect
To verify that the persistence effect seen in Figure 1 was caused by a cosmic ray, we took the uncalibrated ‘uncal’ files for the second eclipse observation and calculated the change in data number between the groups before and after the cosmic ray hit. This occurred between groups 7 and 8 of the 524th integration. To correct for the expected increase in data number due to flux from the target star, thermal background, etc. we subtracted the median frame of all integrations divided by the number of groups. This procedure had minimal effect as this expected increase in data number was far smaller (of the order of hundreds of data numbers) than the size of the jump from the cosmic ray (16,000 data numbers). In Figure 17 we overlay this increase in data number between these two groups with the pixel light curves centred around this cosmic ray event. These pixel light curves are from the reduction described in Section 4.1 which used the jump step correction and also performed an additional custom median outlier clipping step on the stage 1 and 2 processed data. These steps remove the significant jump in flux observed in many of these light curves in the integration containing the cosmic ray hit, but they do not perform any correction for this longer duration persistence artifact.
@@
In summary, persistence effects due to cosmic rays hitting individual pixels may be a widespread phenomenon for which the pixel-fitting method could be beneficial for weighting away from them. If these persistence effects could be modelled then it may be possible to directly fit them in the mean function of the pixel-fitting method (e.g. fitting an exponential ramp to the affected pixels) rather than weight away from them, potentially improving S/N. This could be more important if a strong cosmic ray were to affect one of the central pixels of the PSF or in the case of a high rate of cosmic ray hits as in the eclipse of GJ-357b.
@@
Appendix B Effect of intrapixel sensitivity variations and flat fielding errors
@@
As described in Section 2.1, there are multiple reasons why intrapixel sensitivity should be significantly reduced in MIRI time series compared to Spitzer. To confirm that it can be assumed to be negligble for MIRI, we performed tests to examine to what level intrapixel sensitivity could affect our observations. A related issue is flat fielding, while we did run the flatfield step in the JWST pipeline for the pixel-fitting analyses, any inaccuracies in the flat fielding would result in uncorrected for variations in sensitivity between different pixels (in contrast to intrapixel sensitivity which describes sensitivity variations within a pixel). We can examine both of these effects using simulated observations, similar to what has been performed for Spitzer (Morello 2015).
@@
All of our tests follow the same procedure. First, we generate a possible pixel sensitivity map. As we were unable to find any measurements of intrapixel sensitivity for MIRI in the literature, we tested various possibilities, shown in Figure 18. We considered a ”gradient” map in which the pixel surface smoothly varies in sensitivity with a constant gradient from a relative efficiency of 95% in the top left corner to 105% in the bottom right corner. We also considered a ”quadratic” map where the efficiency varied as a function of the square of the distance from the centre of the pixel. The normalised map has an efficiency ranging from 103% at the pixel centre to 93% at the edges. Our ‘random’ sensitivity map was generated using a GP with a 2D squared-exponential kernel function with correlation a function of the and position within a pixel. The height scale for this GP produced sensitivity variations between 95% to 105% relative efficiency. Each map was generated as a grid of points within each pixel. We then copied this map for a grid of pixels, creating an overall sub-pixel resolution sensitivity map simulating 225 total pixels.
@@
In addition to testing the scenario where all pixels share the same sensitivity map, we also generated maps which used randomised gradient maps, quadratic maps and random 2D GP maps which were independent for each pixel. These maps also exhibit 10% variations in efficiency across each pixel. To test flat fielding errors, we generated sub-pixel maps which had constant efficiency across each surface, but we varied this constant for each pixel with a standard deviation of 1% and centred around 100%. The MIRI reference file for flat fielding exhibits 1% level variations between pixels, so if we run the flatfield step then we are modelling these maps being incorrect with a standard deviation of 1%.
@@
Our results were that all of the intrapixel sensitivity maps and flat fielding errors were completely negligible, never producing greater than a couple ppm of additional noise in our aperture extracted light curves (see Table 9). The pointing variation shows minimal correlation in time so most of this variation behaves like white noise and is tiny compared to the level of photometric noise in our observations (800 ppm). We note that the noise would be added in quadrature to photon noise, so even 40 ppm of noise from intrapixel sensitivity would only increase the total aperture extracted noise by ppm in combination with 800 ppm of photon noise. These effects are therefore more than an order of magnitude below the level that would impact our observations, meaning that significantly more extreme sub-pixel sensitivity variations would be needed to have any noticeable effect. While the maximum effect in a particular pixel could be a couple hundred ppm (see the ‘Max pixel (ppm)’ column in Table 9), this was largely only in the outer regions of the PSF with the lowest S/N and the photon noise in any individual pixel for the real observations is never better than 3000 ppm even in the central pixels, so this effect should also be negligible for the pixel-fitting method. The exact values in Table 9 would vary if different random draws of sensitivity maps were generated, but as the effect is so far below the level of photon noise it appears unlikely this would ever have a significant impact.
It is worth repeating that the Si:As detector technology used for MIRI is expected to show very minor intrapixel sensitivity, so extreme variations across each pixel surface is expected to be highly unlikely. This is similar to what was found with the Si:As detectors on Spitzer, which did not display the same level of intrapixel sensitivity as the shorter wavelength m and m channels which used InSb detectors (Knutson et al. 2008). The narrower PSF widths from the shorter wavelength filters (e.g. F560W) may show slightly more variation, although rerunning the same tests with a reduced FWHM = 1.882px matching that of the F560W filter (and setting the aperture radius to match this FWHM), we still find that all maps produce ppm of noise. This demonstrates that the exquisite pointing stability of JWST is enough to result in minimal intrapixel sensitivity, even if the pixels do demonstrate 10% level variations in sensitivity across their surfaces. Therefore, it appears unlikely that intrapixel sensitivity or flat fielding errors have any significant effect on time series observations with MIRI/Imaging.